Given the following information about a closed oil-rich economy which does not impose any tax on its citizens to finance its expenditure, you are required to derive mathematically and explain using graphs the impact of a change in each of the key exogenous variables on the endogenous variables.
Y = C + I + G
C = C(Y), I = I(r), G = Go, Md = L(Y,r)
1
Expert's answer
2015-10-09T00:00:45-0400
If a closed oil-rich economy does not impose any tax on its citizens to finance its expenditure, Y = C + I + G, C = C(Y), I = I(r), G = Go, Md = L(Y,r), then the increase in exogenous variables as C, I, G and L will cause the endogenous variable Y and Md to increase. Increase in such variable as r will cause Md to decrease.